1) A 6.5 kg ball is launched upwards from the ground with a speed of 21 m/s.

a) How much kinetic energy does the ball have?

b) If all of the ball's kinetic energy is converted to gravitational potential energy, how high will the ball go?



1
Expert's answer
2021-07-14T10:15:51-0400

(a)

"KE=\dfrac{1}{2}mv^2,""KE=\dfrac{1}{2}\cdot6.5\ kg\cdot(21\ \dfrac{m}{s})^2=1433.25\ J."

(b)

"KE=GPE=mgh,""h=\dfrac{KE}{mg}=\dfrac{1433.25\ J}{6.5\ kg\cdot9.8\ \dfrac{m}{s^2}}=22.5\ m."